OpenStudy (anonymous):

HELPPPPP !!! Add and simplify: x+1/x + x-3/3x

jimthompson5910 (jim_thompson5910):

the expression is \[\large \frac{x+1}{x} + \frac{x-3}{3x}\] right?

OpenStudy (anonymous):

yeah @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{x+1}{x} + \frac{x-3}{3x}\] \[\large \frac{3(x+1)}{3x} + \frac{x-3}{3x}\] \[\large \frac{3x+3}{3x} + \frac{x-3}{3x}\] \[\large \frac{3x+3+x-3}{3x}\] \[\large \frac{4x+0}{3x}\] \[\large \frac{4x}{3x}\] \[\large \frac{4}{3}\] --------------------------- So, \[\large \frac{x+1}{x} + \frac{x-3}{3x}\] simplifies to \[\large \frac{4}{3}\] In other words, \[\large \frac{x+1}{x} + \frac{x-3}{3x} = \frac{4}{3}\] is true for all values of x where \(\large x \neq 0\)

jimthompson5910 (jim_thompson5910):

So before you can add the fractions, you need to get all the denominators equal to the LCD (which in this case is 3x)